OpenStudy (anonymous):

3p-1 = 5(p -1) - 2(7-2p) ?? solve

OpenStudy (anonymous):

3p-1 = 5(p -1) - 2(7-2p) 3p-1=5p-5-14+4p 3p-1=9p-19 -3p -3p -1=6p-19 +19 +19 18=6p p=3

OpenStudy (anonymous):

so first we gotta do is 5(p-1)=which is 5p-5 and do -2(7-2p) which is -14+4p so equation is 3p-1=5p-5-14+4p => 3p-1=9p-19 if ur solving what P is than u have to move 19 to other side and 3p to other side so its going to be 60=18 ad u divide both by 6 and p equals 3